# Magnetic Pseudo-differential Operators with Hörmander Symbols Dominated by Tempered Weights

### v1

@@ 1 Introduction
In this paper we study classical Hörmander symbols with uniform decay/growth properties for all derivatives where the decay/growth is governed by a tempered weight (temperate weight [Nicola2010], order function [Zworski2012]). For these symbol classes we derive a characterization of the matrices for the corresponding magnetic pseudo-differential operators (Theorem 2.4). These matrices have a well-defined matrix product with each other, and consequently, we obtain a symbol calculus (Corollary 3.3). Also, we give criteria on the tempered weight for the corresponding magnetic pseudo-differential operators to be bounded, compact, and in the Schatten-classes (Theorems 4.1, 4.2, and 4.3).
@@ 1.1 Symbol Classes and Magnetic Pseudo-differential Operators
$\|\Phi\|_{S_{0}(M),n}=\sum_{\gamma\in{\mathbb{N}}_{0}^{d},|\gamma|\leq n}\sup_{x,\xi\in{\mathbb{R}}^{d}}M(x,\xi)^{-1}|\partial^{\gamma}\Phi(x,\xi)|$
@@ 2.2 Matrix Representation
    If $(N_{\tilde{a},\tilde{\beta}})_{\tilde{\alpha},\tilde{\beta}\in{\mathbb{Z}}^{2d}}$ are complex numbers such that
@@
    then $\sum_{\tilde{\alpha},\tilde{\beta}\in{\mathbb{Z}}^{2d}}N_{\tilde{a},\tilde{\beta}}(\mathcal{W}_{t}^{A})^{-1}\mathcal{G}_{\tilde{\alpha}}^{A}\otimes\overline{\mathcal{G}_{\tilde{\beta}}^{A}}$ converges uniformly on compacts towards a member of $S_{0}(M)$ and each semi-norms of the limit are bounded by a constant times (2.6) for some $n,m$.
@@
$\left|\partial_{x,y,z}^{\delta}\int_{<x,y,z>}B\right|\leq C\langle x-y\rangle^{|\delta|}\langle y-z\rangle^{|\delta|}$ (2.7)
@@
Proof of 2. We want to prove that $\sum_{\tilde{\alpha},\tilde{\beta}\in{\mathbb{Z}}^{2d}}N_{\tilde{a},\tilde{\beta}}(\mathcal{W}_{t}^{A})^{-1}\mathcal{G}_{\tilde{\alpha}}^{A}\otimes\overline{\mathcal{G}_{\tilde{\beta}}^{A}}$ converges uniformly on compacts towards an element of $S_{0}(M)$. We do this by controlling the sum, and the sum of term-wise derivatives, in the differences $\alpha-\beta,\alpha^{\prime}-\beta^{\prime}$ and sums $t\alpha+(1-t)\beta,(1-t)\alpha^{\prime}+t\beta^{\prime}$.
@@
with $C_{3},C_{4}>0$ using (1.2). This last estimate shows that $\sum_{\tilde{\alpha},\tilde{\beta}\in{\mathbb{Z}}^{2d}}N_{\tilde{a},\tilde{\beta}}(\mathcal{W}_{t}^{A})^{-1}\mathcal{G}_{\tilde{\alpha}}^{A}\otimes\overline{\mathcal{G}_{\tilde{\beta}}^{A}}$ converges uniformly on compacts to a $C^{\infty}({\mathbb{R}}^{2d})$ function, and furthermore, this limit is in $S_{0}(M)$.∎
@@ 3.1 Continuous Change of Quantization
Thirdly, the criterion in (2.6) is independent of $t$, so if $(N_{\tilde{a},\tilde{\beta}})_{\tilde{\alpha},\tilde{\beta}\in{\mathbb{Z}}^{2d}}$ satisfy the criterion for one $t\in{\mathbb{R}}$, it satisfies it for all. To see this, suppose $t,s\in[0,1]$ and $\tilde{\alpha},\tilde{\beta}\in{\mathbb{Z}}^{2d}$. Then from (1.1)
@@
The corollary follows directly from the above. Concerning the continuity of $\mathfrak{L}_{t,s}$, it follows from first applying Theorem 2.4 2 to a semi-norm of $\mathfrak{L}_{t,s}$ and then Theorem 2.4 1. It is also easy to see that $\mathfrak{L}_{t,s}^{-1}=\mathfrak{L}_{s,t}$. ∎
@@ 4.2 Schatten-class Properties
Before the proof we note, as a consequence of (1.1), that $M$ is in $L^{p}({\mathbb{R}}^{2d})$ if and only if the values of $M$ on a lattice is $p$-summable, e.g. $(M(\tilde{\gamma}))_{\tilde{\gamma}\in{\mathbb{Z}}^{2d}}\in\ell^{p}({\mathbb{Z}}^{2d})$.

### v2

@@ 1 Introduction
In this paper we study classical Hörmander symbols with uniform decay/growth properties for all derivatives where the decay/growth is dominated by a tempered weight (temperate weight [Nicola2010], order function [Zworski2012]). For these symbol classes we derive a characterization of the matrices for the corresponding magnetic pseudo-differential operators (Theorem 2.4). These matrices have a well-defined matrix product with each other, and consequently, we obtain a symbol calculus (Corollary 3.3). Also, we give criteria on the tempered weight for the corresponding magnetic pseudo-differential operators to be bounded, compact, and in the Schatten-classes (Theorems 4.1, 4.2, and 4.3).
@@ 1.1 Symbol Classes and Magnetic Pseudo-differential Operators
$\|\Phi\|_{S_{0}(M),n}=\sum_{\gamma\in{\mathbb{N}}_{0}^{2d},|\gamma|\leq n}\sup_{x,\xi\in{\mathbb{R}}^{d}}M(x,\xi)^{-1}|\partial^{\gamma}\Phi(x,\xi)|$
@@ 2.2 Matrix Representation
    If $(N_{\tilde{\alpha},\tilde{\beta}})_{\tilde{\alpha},\tilde{\beta}\in{\mathbb{Z}}^{2d}}$ are complex numbers such that
@@
    then $\sum_{\tilde{\alpha},\tilde{\beta}\in{\mathbb{Z}}^{2d}}N_{\tilde{\alpha},\tilde{\beta}}(\mathcal{W}_{t}^{A})^{-1}\mathcal{G}_{\tilde{\alpha}}^{A}\otimes\overline{\mathcal{G}_{\tilde{\beta}}^{A}}$ converges uniformly on compacts towards a member of $S_{0}(M)$ and each semi-norm of the limit are bounded by a constant times (2.6) for some $n,m$.
@@
$\left|\partial_{x,y,z}^{\delta}\int_{<x,y,z>}B\right|\leq C\langle x-y\rangle\langle y-z\rangle$ (2.7)
@@
Proof of 2. We want to prove that $\sum_{\tilde{\alpha},\tilde{\beta}\in{\mathbb{Z}}^{2d}}N_{\tilde{\alpha},\tilde{\beta}}(\mathcal{W}_{t}^{A})^{-1}\mathcal{G}_{\tilde{\alpha}}^{A}\otimes\overline{\mathcal{G}_{\tilde{\beta}}^{A}}$ converges uniformly on compacts towards an element of $S_{0}(M)$. We do this by controlling the sum, and the sum of term-wise derivatives, in the differences $\alpha-\beta,\alpha^{\prime}-\beta^{\prime}$ and sums $t\alpha+(1-t)\beta,(1-t)\alpha^{\prime}+t\beta^{\prime}$.
@@
with $C_{3},C_{4}>0$ using (1.2). This last estimate shows that $\sum_{\tilde{\alpha},\tilde{\beta}\in{\mathbb{Z}}^{2d}}N_{\tilde{\alpha},\tilde{\beta}}(\mathcal{W}_{t}^{A})^{-1}\mathcal{G}_{\tilde{\alpha}}^{A}\otimes\overline{\mathcal{G}_{\tilde{\beta}}^{A}}$ converges uniformly on compacts to a $C^{\infty}({\mathbb{R}}^{2d})$ function, and furthermore, this limit is in $S_{0}(M)$.∎
@@ 3.1 Continuous Change of Quantization
Thirdly, the criterion in (2.6) is independent of $t$, so if $(N_{\tilde{\alpha},\tilde{\beta}})_{\tilde{\alpha},\tilde{\beta}\in{\mathbb{Z}}^{2d}}$ satisfy the criterion for one $t\in{\mathbb{R}}$, it satisfies it for all. To see this, suppose $t,s\in[0,1]$ and $\tilde{\alpha},\tilde{\beta}\in{\mathbb{Z}}^{2d}$. Then from (1.1)
@@
The corollary follows directly from the above. Concerning the continuity of $\mathfrak{L}_{t,s}$, it follows from first applying Theorem 2.4 2 to a semi-norm of $\mathfrak{L}_{t,s}\Phi$ and then Theorem 2.4 1. It is also easy to see that $\mathfrak{L}_{t,s}^{-1}=\mathfrak{L}_{s,t}$. ∎
@@ 4.2 Schatten-class Properties
Before the proof we note, as a consequence of (1.1), that $M$ is in $L^{p}({\mathbb{R}}^{2d})$ if and only if $M$ is measurable and the values of $M$ on a lattice is $p$-summable, e.g. $(M(\tilde{\gamma}))_{\tilde{\gamma}\in{\mathbb{Z}}^{2d}}\in\ell^{p}({\mathbb{Z}}^{2d})$.
@@
Remark 4.5.
The boundedness, compactness, and Schatten-class properties presented in Theorem 4.1, 4.2, and 4.3, can be extended to the magnetic Sobolev spaces presented in [Iftimie2007] using e.g. the methods in [LeeLein2025, Section 5.2].